### v1

@@ 5 Numerical experiments
In this section, we present several numerical experiments illustrating the potential of our minimal residual rational Krylov subspace method (labeled MR-RKSM in the following) in solving long sequences of shifted linear systems. We compare our solver with two state-of-the-art schemes. The first one is the extended Krylov subspace method presented in [32]***The implementation of this method has been obtained by slighlty modifying the function kpik designed for Lyapunov equations [31] and available at https://www.dm.unibo.it/~simoncin/software.html. Similarly to our solver, this scheme relies on the matrix equation formulation (5) of the problem (1) and it solves it by left projection. On the other hand, the adopted approximation space is the extended Krylov subspace (6) and a Galerkin condition is imposed to compute the solution. In the following we refer to this scheme as G-EKSM. The second one is restarted FOM [30]. We preferred this method over restarted GMRES [11] as we will consider difficult problems with nonsymmetric matrices $A$ and possibly complex shifts $s_{j}$ where restarting GMRES may not be possible. This is not an issue for FOM where the residuals are always collinear by construction and restarting is always an option. In the following, we will denote restarted FOM as FOM(100) since we will restart it every 100 iterations, if necessary. Moreover, we set the maximum number of restarting cycles equal to 10.
@@
All the experiments reported in this paper have been run using Matlab (version 2024b) on a machine with a 1.2GHz Intel quad-core i7 processor with 16GB RAM and an Ubuntu 20.04.2 LTS operating system.
We would like to mention that during our numerical testing, we examined the performance of the GCRO-DR method proposed in [21]; see also [22, 24] for a Matlab implementation and a thorough discussion on certain computational aspects of this algorithm. This method is designed for the solution of general parametrized linear systems. Even though (1) can be seen as an instance of this large class of problems, the GCRO-DR implementation available in [23] is not well-suited for a scenario where all the shifts are given. Indeed, in this implementation one is forced to solve the linear systems in (1) one at a time, in a sequential manner. This drawback is overcome in the implementation [34]. On the other hand, this routine builds a block polynomial Krylov subspace for (5) with the whole $b\mathbf{1}_{\ell}^{T}$ as initial block, ignoring its low rank. This issue led to Out-of-Memory errors in our experiments as we always use large values of $\ell$. For these reasons, no results achieved by GCRO-DR are reported.
@@
on $[0,1]^{d}$ with zero Dirchlet boundary conditions. In the expression above, $\nu>0$ is the viscosity parameter and $w$ is the convection vector. The vector $b\in\mathbb{R}^{n^{d}}$ in (1) has random entries drawn from the normal distribution.
We start by setting $d=2$, $n=100$, so that $A$ is 10 000×10 $\times 10\,000$ 10 000, $\nu=0.5$, and $w=(3y(1-x^{2}),-2x(1-y^{2}))$. We consider three sets of =1 $\ell=1\,000$ shifts, each having an increasing level of difficulty. The first set is composed of only real shifts (real) chosen as $\ell$ logarithmically spaced points between $-10^{6}$ and $-10^{-6}$. The second set is given by complex shifts coming in conjugate pairs (complex - conjugate pairs). In particular, we pick $\ell/2$ logarithmically spaced points $\theta_{j}$ between $-10^{6}$ and $-10^{-6}$ and consider the set of $\ell$ shifts given by all the pairs $\rm i\cdot\theta_{j}$, $-\rm i\cdot\theta_{j}$ where $\rm i$ denotes the imaginary unit. The construction of the third set where we have complex shifts with no conjugate pairs (complex - no conjugate pairs) is a little more involved. In particular, we select $\ell$ shifts such that $s_{j}=c+\rho(\cos(\theta_{j})+v\rm{i}\sin(\theta_{j}))$, where $\theta_{j}=2\pi j/\ell$, $c=-223.81+5\rm i$, $\rho=500$, and $v=1$, for $j=1,\ldots,\ell$.
@@
Table 2: Example 5.1 (2D). Performance of the different methods by varying the nature of the =1 $\ell=1\,000$ shifts $s_{j}$ in (1). The reported timings are in seconds. “*” means no convergence within the prescribed maximum number of iterations (and restarting cycles).
@@
In Table 2, we collect the results achieved by MR-RKSM, G-EKSM, and FOM(100). In particular, we report the overall number of iterations, the rank of the computed solution, and the running time in seconds. For FOM(100) we also report the number of restarting cycles performed by the method. When we report a number of cycles equal to cycle we mean that FOM(100) stopped while performing the cycle-th restart. Note that MR-RKSM and G-EKSM do not perform any restart. Moreover, we adopt $\text{Rank}(X_{m})$ as an indicator of the memory usage of the different methods. In particular, MR-RKSM and G-EKSM represent the solution in low-rank format by allocating $\text{Rank}(X_{m})$ vectors of length $n^{2}$ and other $\text{Rank}(X_{m})$ of length $\ell$. On the other hand, by construction FOM(100) stores all the $\ell$ solutions to (1) as this is necessary to update the current approximation when restarting is applied. This means that FOM(100) is not able to take advantage of the low rank of $X=[x^{(1)},\ldots,x^{(\ell)}]$ to cut its memory demand.
@@
From the results in Table 2 we can see that all three methods perform rather well for the sets of shifts real and complex - conjugate pairs, with MR-RKSM being the fastest and the one with the lowest memory requirements. On the other hand, it is with the shifts complex - no conjugate pairs that the performance of MR-RKSM turns out to be outstanding. Indeed, this is the only method able to converge at all. Moreover, convergence is achieved in a small number of iterations and a rather fast running time. G-EKSM and FOM(100) do not converge in 100 iterations and 10 restarting cycles, respectively. In particular, the relative residual norm attained by G-EKSM exhibits a very oscillating trend in the first 50 iterations and then starts to slowly decrease, settling at $\mathcal{O}(10^{-4})$ at the 100th iteration. We remind the reader that, at this point, the rank of the approximate solution computed by G-EKSM is already 200, much larger than the one achieved by MR-RKSM. On the other hand, the FOM(100) relative residual norm shows this oscillating behavior throughout all iterations and cycles with no significant decrease. This is clearly visible by looking at Figure 1 where we depict the convergence history of the three methods.
@@
Figure 1: Example 2.3 (2D). Relative residual norm achieved by MR-RKSM (blue, solid line), G-EKSM (red, dashed line), and FOM(100) (black, dotted line). For FOM(100) we report the relative residual norm of the first 100 iterations, namely during the first cycle. A similar behavior is also observed in the following restarting cycles.
@@
We now construct a different $A$ by setting $d=3$, $n=50$, so that $A$ is now 125 000×125 $\times 125\,000$ 125 000, and $w=(x\cos(x),y\sin(y),\exp(z^{2}-1))$. The viscosity parameter $\nu$ and the shift sets are as before. It is well-known that iterative solvers work better than their direct counterparts when dealing with linear systems stemming from the discretization of 3D problems. Therefore, we now consider implementations of MR-RKSM and G-EKSM where the linear systems involved in the basis construction are solved by GMRES(50) [28], where we allow for 100 restarting cycles at most, and preconditioned by an incomplete LU factorization with no fill-in. We report the corresponding results in Table 3.
@@
Table 3: Example 5.1 (3D). Performance of the different methods by varying the nature of the =1 $\ell=1\,000$ shifts $s_{j}$ in (1). The reported timings are in seconds. “*” means no convergence within the prescribed maximum number of iterations (and restarting cycles).
The iterative solution of the inner (shifted) linear systems makes the computational gap between the matrix-equation-based solvers (MR-RKSM and G-EKSM) and FOM(100) much narrower, especially for real and complex - conjugate pairs shifts, to the point that FOM(100) turns out to be the fastest method in the case of complex - conjugate pairs shifts. On the other hand, MR-RKSM and G-EKSM are able to capitalize on the low-rank structure of the solution and remarkably reduce the storage demand. As for the 2D example, also in this case MR-RKSM is the only method able to converge when adopting complex - no conjugate pairs shifts. The quite high computational time needed by MR-RKSM in this case is due to GMRES(50). Indeed, for some poles $\xi_{j}$, GMRES(50) needs a large number of restarting cycles for solving the related shifted linear systems with $A+\xi_{j}I$. This can be possibly fixed by employing more performing preconditioning operators.
@@
[…]
Also for this example, G-EKSM and FOM(100) do not perform well due to the nature of the shifts. We thus decided to skip them here. On the other hand, in addition to testing MR-RKSM, we document the running time devoted to sequentially solving all the shifted linear systems in (1) by “backslash”. Indeed, due to the small value of $n$, one may think that the latter is a more suitable option. Nevertheless, we are going to show how our low-rank approach is competitive even in this scenario. Moreover, we compare the pole selection strategy we proposed in section 3.1 with state-of-the-art procedures agnostic to the underlying shifted linear systems problem. In particular, we employ the strategy called ADM in [6] and implemented in the function newpole_adaptive_ADM.m available in the package rk_adaptive_sylvesterAvailable at https://github.com/numpi/rk_adaptive_sylvester/tree/main. Notice, however, that, similarly to any other state-of-the-art scheme for pole selection in rational Krylov methods for Sylvester equations, the strategies presented in [6] are based on an expression of the residual matrix obtained by imposing a Galerkin condition rather than a minimal residual one as we do in our approach. Therefore, the findings from [6] may need some adjustments to fully adapt to our framework. This is clearly beyond the scope of this paper. Our goal here is just to illustrate how the simple and somehow natural pole selection strategy from section 3.1 works better than off-the-shelf, involved alternatives that are available in the literature.
@@
MR-RKSM (Poles as in sec. 3.1) 42 42 1.58
MR-RKSM (ADM poles [6]) 88 88 4.87
backslash - - 5.41
@@
MR-RKSM (Poles as in sec. 3.1) 41 41 1.69
MR-RKSM (ADM poles [6]) 87 87 6.81
backslash - - 10.23
@@
MR-RKSM (Poles as in sec. 3.1) 42 42 2.57
MR-RKSM (ADM poles [6]) 87 87 11.88
backslash - - 21.38
@@
In Table 4, we report the results obtained by changing the number $\ell$ of shifts. The reported timings showcase the potential of our rational Krylov approach even when dealing with medium-sized matrices $A$ and a moderate number of shifts $\ell$. Moreover, the timings of the solution process based on backslash clearly grow linearly with $\ell$ whereas MR-RKSM (with our pole selection) is not really affected by this drawback. The poles computed by the ADM strategy from [6] are not as effective as the ones computed by following section 3.1. Indeed, MR-RKSM with the latter poles performs way fewer iterations (about 41 vs 87) than what happens by computing the poles as in [6]. It is worth noting that the overall running time of MR-RKSM shows a much more important correlation with $\ell$ when ADM is employed, in contrast to what happens when the strategy from section 3.1 is adopted. This is caused by the cost of ADM which now depends on $\ell$ due to the lack of a right projection in our framework.
@@ 6 Conclusion
We showed once again that a matrix equation formulation is the way to go when dealing with sequences of shifted linear systems of the form (1). This point of view was first explored in [32] where an extended Krylov subspace method equipped with a Galerkin condition was proposed. This paper makes an important step further. In particular, thanks to our minimal residual rational Krylov subspace method, we are now able to efficiently solve difficult problems, with nonsymmetric coefficient matrices $A$ and complex shifts not coming in conjugate pairs. Like any projection technique for matrix equation, our solver works well if and only if the exact solution presents a fast decay in its singular values. We portrayed scenarios where we expect this to happen. Moreover, we generalized the approach to the block case, where we have a number of different right-hand sides $b_{j}$, and proposed a cheap and effective pole selection strategy for the construction of the rational Krylov subspace.
We would like to mention that the minimal residual condition adopted in this paper can be easily integrated in an extended Krylov subspace method by following, e.g., what has been done in [5] in the context of shifted linear systems within the ADI solver for Lyapunov equations. Moreover, even though in our numerical experiments the constructed subspace has always been of very moderate dimension, one could try to adapt the compress-and-restart paradigm presented in [18] for polynomial Krylov subspaces to rational ones. This undergoes the understanding of collinearity of residuals in a matrix equation setting, a concept that, to the best of our knowledge, has never been explored in the literature.

### v2

@@ 5 Numerical experiments
In this section, we present several numerical experiments illustrating the potential of our minimal residual rational Krylov subspace method (labeled MR-RKSM in the following) in solving long sequences of shifted linear systems. We compare our solver with two state-of-the-art schemes. The first one is the extended Krylov subspace method presented in [Simoncini2010]The implementation of this method has been obtained by slighlty modifying the function kpik designed for Lyapunov equations [Sim07] and available at https://www.dm.unibo.it/~simoncin/software.html. Similarly to our solver, this scheme relies on the matrix equation formulation (5) of the problem (1) and it solves it by left projection. Compared to our proposed method, there are two main differences: (i) the adopted approximation space is the extended Krylov subspace (6) and (ii) a Galerkin condition is imposed to compute the solution. In the following we refer to this scheme as G-EKSM. The second one is restarted FOM [FOM]. We preferred this method over restarted GMRES [RestartedGMRES] as we will consider difficult problems with nonsymmetric matrices $A$ and possibly complex shifts $s_{j}$ where restarting GMRES may not be possible. This is not an issue for FOM where the residuals are always collinear by construction and restarting is always an option. In the following, we will denote restarted FOM as FOM(100) since we will restart it every 100 iterations, if necessary. Moreover, we set the maximum number of restarting cycles equal to 10.
@@
All the experiments reported in this paper have been run using Matlab (version 2024b) on a machine with a 1.2GHz Intel quad-core i7 processor with 16GB RAM and an Ubuntu 20.04.2 LTS operating system. The Matlab code for reproducing the experiments that follow is available at https://github.com/palittaUniBO/MR_RKSM/tree/main.
We would like to mention that during our numerical testing, we examined the performance of the GCRO-DR method proposed in [Parks2006]; see also [Parks2016a, Parks2016] for a Matlab implementation and a thorough discussion on certain computational aspects of this algorithm. This method is designed for the solution of general parametrized linear systems. Even though (1) can be seen as an instance of this large class of problems, the GCRO-DR implementation available in [Parks2016a] is not well-suited for a scenario where all the shifts are given. Indeed, in this implementation one is forced to solve the linear systems in (1) one at a time, in a sequential manner. This drawback is overcome in the implementation [soodhalter_2016_56157]. However, this routine builds a block polynomial Krylov subspace for (5) with the whole $b\mathbf{1}_{\ell}^{T}$ as initial block, ignoring its low rank. This issue led to Out-of-Memory errors in our experiments as we always use large values of $\ell$. For these reasons, no results achieved by GCRO-DR are reported.
@@
on $0,1]^{d}$ with zero Dirchlet boundary conditions. In the expression above, $\nu>0$ is the viscosity parameter and $w$ is the convection vector. The vector $b\in\mathbb{R}^{n^{d}}$ in (1) has random entries drawn from the normal distributionrng(’default’) and unit norm.
We start by setting $d=2$, $n=100$, so that $A$ is 10 000×10 $00\times 10\,00$0, $\nu=0.5$, and $w=(3y(1-x^{2}),-2x(1-y^{2}))$. We consider three sets of $\ell=1\,000$ shifts, each having an increasing level of difficulty. The first set is composed of only real shifts (real) chosen as $\ell$ logarithmically spaced points between $10^{6}$ and $10^{-6}$. The second set is given by complex shifts coming in conjugate pairs (complex - conjugate pairs). In particular, we pick $\ell/2$ logarithmically spaced points $\theta_{j}$ between $10^{6}$ and $10^{-6}$ and consider the set of $\ell$ shifts given by all the pairs $\rm i\cdot\theta_{j}$, $\rm i\cdot\theta_{j}$ where $\rm i$ denotes the imaginary unit. The construction of the third set where we have complex shifts with no conjugate pairs (complex - no conjugate pairs) is a little more involved. In particular, we select $\ell$ shifts such that $s_{j}=c+\rho(\cos(\theta_{j})+v\rm{i}\sin(\theta_{j}))$, where $\theta_{j}=2\pi j/\ell$, $c=-223.81+5\rm i$, $\rho=500$, and $v=1$, for $j=1,\ldots,\ell$.
@@
Table 3: Example 5.1 (2D). Performance of the different methods by varying the nature of the $\ell=1\,000$ shifts $s_{j}$ in (1). The reported timings are in seconds. “*” means no convergence within the prescribed maximum number of iterations (and restarting cycles).
@@
In Table 3, we collect the results achieved by MR-RKSM, G-EKSM, and FOM(100). In particular, we report the overall number of iterations, the rank of the computed solution, and the running time in seconds. For FOM(100) we also report the number of restarting cycles performed by the method. When we report a number of cycles equal to cycle we mean that FOM(100) stopped while performing the cycle-th restart. Note that MR-RKSM and G-EKSM do not perform any restart. Moreover, we adopt $\text{Rank}(X_{m})$ as an indicator of the memory usage of the different methods. In particular, MR-RKSM and G-EKSM represent the solution in low-rank format by allocating $\text{Rank}(X_{m})$ vectors of length $n^{2}$ and other $\text{Rank}(X_{m})$ vectors of length $\ell$. Note that FOM(100) stores all the $\ell$ solutions to (1) as this is necessary to update the current approximation when restarting is applied. This means that FOM(100) is not able to take advantage of the low rank of $X=[x^{(1)},\ldots,x^{(\ell)}]$ to cut its memory demand.
@@
From the results in Table 3 we can see that all three methods perform rather well for the sets of shifts real and complex - conjugate pairs, with MR-RKSM being the fastest and the one with the lowest memory requirements. It is with the shifts complex - no conjugate pairs that the performance of MR-RKSM turns out to be outstanding. Indeed, this is the only method able to converge at all. Moreover, convergence is achieved in a small number of iterations and a rather fast running time. G-EKSM and FOM(100) do not converge in 100 iterations and 10 restarting cycles, respectively. In particular, the relative residual norm attained by G-EKSM exhibits a very oscillating trend in the first 50 iterations and then starts to slowly decrease, settling at $\mathcal{O}(10^{-4})$ at the 100th iteration. We remind the reader that, at this point, the rank of the approximate solution computed by G-EKSM is already 200, much larger than the one achieved by MR-RKSM. The FOM(100) relative residual norm shows oscillating behavior throughout all iterations and cycles with no significant decrease. This is clearly visible by looking at Figure 1 where we depict the convergence history of the three methods. Even though we do not have a full understanding of this phenomenon, the superiority of the rational Krylov subspace over its extended and polynomial counterparts when dealing with shifted linear systems with complex shifts was already documented in [Poles2, Section 3.1] in the context of transfer function approximation. Similarly, it is well-known that polynomial Krylov subspace methods struggle to deal with problems stemming from the discretization of Helmholtz equations [Ernst2012]; a problem related to the solution of shifted linear systems. In any case, further study on the convergence properties of our scheme is necessary to get a complete picture.
@@
Figure 1: Example 5.1 (2D) - shifts: complex - no conjugate pairs. Relative residual norm achieved by MR-RKSM (blue, solid line), G-EKSM (red, dashed line), and FOM(100) (black, dotted line). For FOM(100) we report the relative residual norm of the first 100 iterations, namely during the first cycle. A similar behavior is also observed in the following restarting cycles.
@@
We now construct a different $A$ by setting $d=3$, $n=50$, so that $A$ is now 125 000×125 $00\times 125\,00$0, $\nu=1$, and $w=(x\cos(x),y\sin(y),\exp(z^{2}-1))$. The shift sets are as before. It is well-known that iterative solvers work better than their direct counterparts when dealing with linear systems stemming from the discretization of 3D problems. Therefore, we now consider implementations of MR-RKSM and G-EKSM where the linear systems involved in the basis construction are solved by GMRES(50) [Schultz1986], where we allow for 100 restarting cycles at most, and preconditioned by an incomplete LU factorization with no fill-in. GMRES(50) is stopped as soon as the computed relative residual norm drops below $10^{-9}$, namely one order of magnitude smaller than the outer tolerance. Notice, however, that this inner threshold may be relaxed, i.e., increased, as the MR-RKSM iterations proceed by adapting the inexact rational Krylov approach presented in [inexactRational] for the solution of Lyapunov equations. While this approach would make the results in Proposition 3.4 not holding anymore, the computed rational Krylov space may still be a good approximation space.
We report the results for a fixed inner tolerance in Table 4.
@@
Table 4: Example 5.1 (3D). Performance of the different methods by varying the nature of the $\ell=1\,000$ shifts $s_{j}$ in (1). The reported timings are in seconds. “*” means no convergence within the prescribed maximum number of iterations (and restarting cycles).
The iterative solution of the inner (shifted) linear systems makes the computational gap between the matrix-equation-based solvers (MR-RKSM and G-EKSM) and FOM(100) much narrower, especially for real and complex - conjugate pairs shifts, to the point that FOM(100) turns out to be the fastest method in the case of complex - conjugate pairs shifts. Nevertheless, MR-RKSM and G-EKSM are able to capitalize on the low-rank structure of the solution and remarkably reduce the storage demand. As for the 2D example, also in this case MR-RKSM is the only method able to converge when adopting complex - no conjugate pairs shifts. The quite high computational time needed by MR-RKSM in this case is due to GMRES(50). Indeed, for some poles $\xi_{j}$, GMRES(50) needs a large number of restarting cycles for solving the related shifted linear systems with $A+\xi_{j}I$. This can be possibly fixed by employing more performing preconditioning operators.
@@
In the second experiment, we consider as $A$ the matrix qc2534§§§Available at https://math.nist.gov/MatrixMarket/data/NEP/h2plus/h2plus.html which is part of the data set of matrices used for testing eigensolvers in the SLEPc package [SLEPc]. This matrix is of moderate dimension, $A\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{n\times n}$, n=2 534n=2\,534, $b\in\mathbb{R}^{n}$ is again a Gaussian vectorrng(’default’)with unit norm. We consider $\ell$ shifts of the type complex - no conjugate pairs constructed as before. In particular, $s_{j}=c+\rho(\cos(\theta_{j})+v\rm{i}\sin(\theta_{j}))$ where $\theta_{j}=2\pi j/\ell$, $c=-0.8-0.07\rm i$, $\rho=0.2$, and $v=0.1$, for $j=1,\ldots,\ell$ and different values of $\ell$.
Also for this example, G-EKSM and FOM(100) do not perform well due to the nature of the shifts. We thus decided to skip them here. In addition to testing MR-RKSM, we document the running time devoted to sequentially solving all the shifted linear systems in (1) by “backslash”. Indeed, due to the small value of $n$, one may think that the latter is a more suitable option. Nevertheless, we are going to show how our low-rank approach is competitive even in this scenario. Moreover, we compare the pole selection strategy we proposed in section 3.1 with state-of-the-art procedures agnostic to the underlying shifted linear systems problem. In particular, we employ the strategy called ADM in [Poles1]. To this end, we slightly modify the function rk_adaptive_sylvester.m, available in the package rk_adaptive_sylvesterAvailable at https://github.com/numpi/rk_adaptive_sylvester/tree/main, designed for general Sylvester equations of the form $AX-XB-uv^{T}=0$, in order to perform only a left projection and solve (16) at the projected level. Notice, however, that, similarly to any other state-of-the-art scheme for pole selection in rational Krylov methods for Sylvester equations, the strategies presented in [Poles1] are based on an expression of the residual matrix obtained by imposing a Galerkin condition rather than a minimal residual one as we do in our approach. Therefore, the findings from [Poles1] may need some adjustments to fully adapt to our framework. This is clearly beyond the scope of this paper. Our goal here is just to illustrate how the simple and somehow natural pole selection strategy from section 3.1 works better than off-the-shelf, involved alternatives that are available in the literature.
@@
MR-RKSM (Poles as in sec. 3.1) 38 39 0.78
MR-RKSM (ADM poles [Poles1]) 127 128 7.84
backslash - - 2.92
@@
MR-RKSM (Poles as in sec. 3.1) 38 39 1.05
MR-RKSM (ADM poles [Poles1]) 127 128 11.80
backslash - - 6.40
@@
MR-RKSM (Poles as in sec. 3.1) 38 39 1.60
MR-RKSM (ADM poles [Poles1]) 127 128 20.27
backslash - - 12.57
@@
In Table 5, we report the results obtained by changing the number $\ell$ of shifts. The reported timings showcase the potential of our rational Krylov approach even when dealing with medium-sized matrices $A$ and a moderate number of shifts $\ell$. Moreover, the timings of the solution process based on backslash clearly grow linearly with $\ell$ whereas MR-RKSM (with our pole selection) is not really affected by this drawback. The poles computed by the ADM strategy from [Poles1] are not as effective as the ones computed by following section 3.1, to the point that MR-RKSM with ADM poles is less performing than the naive backslash solution. In Figure 2, for $\ell=1024$, we report $\|r_{m}^{(j)}\|$ in logarithmic scale for all the iterations $m$ ($x$-axis) and shift index $j$ ($y$-axis) in case of our pole selection strategy (left) and with the ADM poles (right). We can notice a very different behavior. In particular, in Figure 2 (left) we can spot several blue horizontal lines. Those are the residual norms related to those shifted linear systems whose shift is used as rational Krylov pole. Indeed, as shown in Proposition 3.4, those residual norms become zero as soon as this happens. This is not the case for the ADM poles that bring the residual norms down to the target accuracy much more slowly. It is also worth noting that the overall running time of MR-RKSM shows a much more important correlation with $\ell$ when ADM is employed, in contrast to what happens when the strategy from section 3.1 is adopted. This is caused by the cost of ADM which now depends on $\ell$ due to the lack of a right projection in our framework.
Figure 2: Example 5.2, $\ell=1024$. Heatmap where we report $\|r_{m}^{(j)}\|$ in logarithmic scale for all iterations $m$ ($x$-axis) and shift indeces $j$ ($y$-axis). Left: MR-RKSM with poles as in sec. 3.1. Right: MR-RKSM with ADM poles [Poles1].
@@ 6 Conclusion
We showed once again that a matrix equation formulation is the way to go when dealing with sequences of shifted linear systems of the form (1). This point of view was first explored in [Simoncini2010] where an extended Krylov subspace method equipped with a Galerkin condition was proposed. This paper makes an important step further. In particular, thanks to our minimal residual rational Krylov subspace method, we are now able to efficiently solve difficult problems, with nonsymmetric coefficient matrices $A$ and complex shifts not coming in conjugate pairs. Like any projection technique for matrix equation, our solver works well if and only if the exact solution presents a fast decay in its singular values. We portrayed scenarios where we expect this to happen. Moreover, we generalized the approach to the block case, where we have a number of different right-hand sides $b^{(j)}$, and proposed a cheap and effective pole selection strategy for the construction of the rational Krylov subspace.
We would like to mention that the minimal residual condition adopted in this paper can be easily integrated in an extended Krylov subspace method by following, e.g., what has been done in [Benneretal2023] in the context of shifted linear systems within the ADI solver for Lyapunov equations. Moreover, even though in our numerical experiments the constructed subspace has always been of very moderate dimension, one could try to adapt the compress-and-restart paradigm presented in [Kressneretal2021] for polynomial Krylov subspaces to rational ones. This requires an understanding of the collinearity properties of residuals in matrix equations, which, to the best of our knowledge, has never been explored in the literature.
We also observed that each of the different Krylov approaches may be preferable for a particular class of shifted linear systems; however, identifying the most efficient method is not straightforward in the absence of spectral information about the matrices involved. Although polynomial Krylov methods do not require the solution of linear systems, they often exhibit slow convergence. Moreover, the large number of iterations typically needed for convergence limits the extent to which the low-rank structure of the solution can be effectively exploited.
Extended Krylov methods, on the other hand, involve solving a sequence of linear systems that share the same coefficient matrix, which can be advantageous. Nevertheless, we observed that for certain examples motivated by e.g., the contour integral method for eigenvalue problem, the convergence of these methods may deteriorate, necessitating a large number of iterations and thereby diminishing the benefits of having a low-rank solution.
Our proposed approach which is based on rational Krylov methods requires the solution of a number of shifted linear systems, which is feasible when an efficient linear solver is available. The principal advantage of this method is its rapid convergence combined with very low memory requirements. Indeed, our numerical experiments indicate that the memory usage is close to optimal.